i have 216 parts which are bad and 15 parts that are good. what is the percentage of good parts?
Rustybutt

Found 2 solutions by Cromlix, stanbon:
Answer by Cromlix(4381)

Hi there,
In total you have 216 + 15 = 231
15/231 x 100 = 6.5%
Hope this helps :-)

Answer by stanbon(75887)

i have 216 parts which are bad and 15 parts that are good. what is the percentage of good parts?
------
15/(216+15) = 0.0649 = 6.49%
-------------
Cheers,
Stan H.
--------